Solve each system by the substitution method. Determine whether the equations are independent, dependent, or inconsistent.
29.) 2x + y = 9
2x – 5y = 15

The books answer is this what I don't understand is how the book came out with this answer. Can you help me solve this please?
{(5, -1)}, independent
Thank you-

Answer by longjonsilver(2297)

substitution method...

eqn1 can be rewritten as 2x = 9-y. Now put this into the second equation as:

9-y - 5y = 15 and now simplfy
9 - 6y = 15
-6y = 6
y = 6/-6
--> y = -1

sub this back into 2x = 9-y

2x = 9-(-1)
2x = 9+1
2x = 10
x = 10/2
--> x = 5

jon.